

********************************************************************
UARS WEEKLY STATUS REPORT - 06 AUG 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
64% of each day's data amounting to about 589 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2003365 has been reprocessed and archived at the GSFC DAAC. Processing
of HEPSA data is temporarily suspended waiting for additional UARS OA data
processing.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004365. Processing of MEPS data is
temporarily suspended waiting for additional UARS OA data processing.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems. The MPG attended the UARS Monthly Status
meeting, held on August 03 to discuss significant events for the
month of July.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 July 2004 (Orbit 70413, GMT Day 206) through
30 July 2004 (Orbit 70518, GMT Day 212).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 206/1301 70421 POWER ON
HRDI 209/1502 70467 TO LOW POWER MODE
PEM 211/1347 70497 AXIS 1 & 2 ON
The beta angle ranged between the angles of 22.2 degrees to 36.8
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 39 degrees on 01 August 2004 (DOY 214).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 1.28 degrees C. Additional
battery performance data for each battery on 23 July and 30 July 2004
are provided in Appendix D.
The clock error ranged from 10.4 to -12.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
HALOE = 02
HRDI = 01
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued testing the new version of the Eclipse ground system
software. There are several DR fixes incorporated in the build, so
testing will take some time. The goal for the testing to be complete
is the end of the July. If all testing is successful the new build
will be released to operations.
The new CMS code to allow for SMA events continued to undergo testing
this week. Problems receiving products from FDF, transferring the
schedule from UPS to CMS, and system drive issues hindered progress.
While the FDF product and system drive issues have been resolved, the
transfer issue is still under investigation.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 August 2004 (DOY 231).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 July 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70421 206/1301 04-217 HALOE POWER ON (PART 1)
70422 206/1441 04-216 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70423 206/1620 04-218 HALOE POWER ON (COMPLETE)
70439 207/1802 04-219 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70467 209/1502 04-220 HRDI TO LOW POWER MODE DUE TO S/C POWER
70486 210/2046 04-221 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70497 211/1347 04-222 PEM AXIS 1 & 2 ON
70504 212/0223 04-223 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70508 212/0839 04-224 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 July 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 July 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70422 206/1443 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70439 207/1809 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70486 210/2102 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70504 212/0225 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70508 212/0846 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 July 2004
23 July 2004 - GMT Day 205 Beta = 04.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-44.8 +4.19/+3.23 +27.2/-12.8 26.2 1.01 18.5
3 +11.2/-72.8 +2.92/+2.29 +25.6/-12.8 26.2 1.00 18.1
30 July 2004 - GMT Day 212 Beta = 36.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-184.8 +5.15/+4.51 +26.4/-14.4 25.6 1.02 20.0
3 +33.6/-072.8 +4.51/+3.23 +25.6/-14.4 25.6 1.02 19.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 July 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT 70477 210/0703 30832 Generic Late Acquisition
04-039 FOT 70492 211/0557 30832 Generic Late Acquisition
04-049 FOT 70501 211/2042 36507 Data Dropout - weather W/S
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-047 FOT 70367 202/2252 N/A WS-4 not responding properly
04-048 FOT 70395 204/1920 36444 Data Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT See List July 04 30832 Generic Late Acquisition
04-049 FOT 70501 211/2042 36507 Data Dropout - weather W/S